# An integrated widely tunable linear isolator based on electro-optic Autler-Townes splitting

### v1

@@ Results
Figure 4: Sideband analysis during isolator operation. (a) The measured first-order sidebands as a function of optical detuning, normalized to the input optical carrier, for the case shown in Fig. 3c. (b) These beat spectra obtained from the heterodyne detection system reveal the carrier and its sidebands (see Supplementary Data Fig. 1), normalized to the input optical carrier. Measurements are performed with a $\Omega_{r}=100$ MHz frequency shift applied to the reference arm of the interferometer; thus $\Omega_{r}$ measures the carrier, $\Omega\pm\Omega_{r}$ correspond to the first-order Stokes and anti-Stokes sideband respectively, and $2\Omega\pm\Omega_{r}$ correspond to the second-order Stokes and anti-Stokes sideband respectively. The anti-Stokes sidebands are notably absent, and beat notes at $\Omega,2\Omega$ correspond to parasitics in the measurement system. (c) We plot the normalized carrier transmission and the Stokes sideband power in forward direction, normalized to input optical carrier, as a function of RF voltage at $\Delta$ = 0. The solid line is a prediction from coupled mode theory [sohn_time-reversal_2018, sohn_electrically_2021] using the experimentally measured device parameters.
Demonstration of electro-optic isolation – Fig. 3 presents an example experiment on the isolator. First, we measured optical transmission through the waveguide in the telecom band, around 1550 nm, which reveals two distinct optical mode families of the resonator (Fig. 3a) with separation of a few GHz. The modes can be identified as TEeven and TEodd based on both coupling to the waveguide and their dispersive effect supported through finite element simulations. Next, we experimentally measure the through-waveguide transmission using a heterodyne detection system (see Methods) which enables resolved measurements on the carrier transmission, as well as any sideband signals that are generated. In this experiment we chose an optical mode pair near 193.98 THz (1545.50 nm) as identified in Fig. 3a with frequency gap of approximately 4.8 GHz and individual loss rates $\kappa_{\textrm{even}}=0.28$ GHz and $\kappa_{\textrm{odd}}=$ 0.53 GHz. We accordingly apply a 3-phase RF stimulus at $\Omega=4.8$ GHz to the electrodes. Fig. 3b presents the evolution of the optical transmission through waveguide as a function of RF voltage (i.e. sinusoidal amplitude of RF stimulus, VRF) for both forward and backward directions. Here, we have we arbitrarily set the TEodd mode as the zero detuning reference point ($\Delta=0$) for the optical frequency axis. In the absence of any applied RF, the transmission function remains identical in both directions. However, with increasing RF voltage, the p-ATS modal hybridization appears for one direction of propagation only. In Fig. 3c we present a superimposed result at the highest applied RF amplitude of 15.88 V, reaching a peak isolation contrast of 47.7 dB with a forward insertion loss of 1.45 dB. This implies IC/IL ratio of $\approx 32.9$ dB/dB which makes the performance comparable to commercial magneto-optic isolators. By fitting to coupled mode theory [sohn_electrically_2021], we determine that this driving condition corresponds to an EO coupling rate $G_{eo}\approx 0.89$ GHz. We measure a 10 dB IC bandwidth of $\approx 150$ MHz and 20 dB IC bandwidth of $\approx 40$ MHz. We tested this device with an estimated in-waveguide optical power up to 5 mW.
While the EO stimulus on this device has an overall traveling wave characteristic, the propagating momentum is generated only synthetically. We confirm this using a RF S11 reflection measurement at the stimulus electrodes (Supplementary Data Figure 2) which confirms their uniformly high reflection and open circuit (capacitive) behavior [orsel_electro-optic_2023]. The residual loss arises from propagation in the measurement cables. For the specific experiment in Fig. 3, the bench-top signal generator power was set at 28 dBm (630 mW), and the RF S11 measurement at 4.8 GHz informs us of a fundamental upper limit of only 12.9 mW (11.09 dBm) power consumed by the device. With improved RF design and a custom source, the power consumption can be reduced significantly. Thus, essentially all of the signal generator voltage (15.88 V) appears directly on the stimulus electrodes due to the open circuit.
Sideband analysis – During experiments we also monitored the anti-Stokes and Stokes sidebands since they are a concern in non-reciprocal devices that use strong spatio-temporal modulation [doerr_optical_2011, doerr_silicon_2014, yu_integrated_2023, gao_thin-film_2024]. Fig. 4a presents measurements of the first-order sidebands as a function of optical detuning $\Delta$, under the operational conditions of Fig. 3c. Naturally, the sideband generation in the forward (phase-matched) direction is shaped by the optical DoS as modified by the p-ATS. In the limit of very strong coupling, the p-ATS dressed states will be fully resolved [kim_complete_2017], and at $\Delta=0$ the optical DoS will reach zero causing the forward direction first-order sideband to be eliminated. In the backward (not phase-matched) direction the sidebands are significantly smaller.
Fig. 4b analyzes the higher-order sidebands in the optical spectrum using the heterodyne interferometer (see Methods and Supplementary Data Figure 1) with the input laser centered on the TEodd mode ($\Delta=0$) and reference arm frequency shift of $\Omega_{r}=100$ MHz. We expect that only first-order Stokes scattering will be notable since only scattering to TEeven can take place. The interferometer spectrum shows the carrier (at $\Omega_{r}$), first Stokes (at $\Omega+\Omega_{r}$), and a much smaller second-order Stokes (at $2\Omega+\Omega_{r}$) signal. We normalize the forward and backward measurements relative to the carrier transmission in the forward direction. These measurements confirm that only the first Stokes sideband is of significance. The first anti-Stokes ($\Omega-\Omega_{r}$) and second anti-Stokes ($2\Omega-\Omega_{r}$) are absent, or below the noise floor, as there are no optical states available within the resonator.
These sideband generation behaviors are readily predictable from coupled mode theory [sohn_electrically_2021, kim_complete_2017] and two regimes of operation have been identified. When the applied stimulus $G_{eo}$ is small, the device acts as a non-reciprocal frequency shifting modulator [sohn_time-reversal_2018]. However, once $G_{eo}>\sqrt{\kappa_{\textrm{even}}\cdot\kappa_{\textrm{odd}}}$ the device enters the strong coupling or p-ATS regime where linear isolation becomes available. With increasing drive level the p-ATS dressed states get resolved further and the ideality of the device monotonically improves as sideband generation reduces monotonically. This behavior and the close match between experiments and theory is confirmed in Fig. 4c.
{adjustwidth*}
-1in-1in Refer to caption
Figure 5: Frequency tuning of the isolator band. (a) The isolator is tested across a 1 THz band (8 nm wavelength range between 1542 – 1550 nm), with isolation contrast and insertion loss measured for 27 distinct mode pairs. The RF stimulus frequency $\Omega$ is set individually in each experiment to match the frequency gap for the mode pair, while VRF is fixed to 15.88 V. All measurements maintain $41-52$ dB contrast across this spectral range and show insertion loss $1.4-2.0$ dB. The $\color[rgb]{1,0,0}\bm{\star}$ indicates the mode pair examined in Figs. 3 and 4. (b) The EO tuner (shown in Fig. 2e) enables fine-tuning, as it produces similar tunability for both optical modes and the frequency gap evolves minimally. Here we present an example meaurement on a different device with mode pair spacing around 11.5 GHz.
@@
Tunability across a wide spectrum – We additionally confirmed the tuning capability of this EO isolator over the range 193.35 – 194.45 THz (approximately 8 nm from 1542 – 1550 nm). For all the mode pairs shown in Fig. 5a, we identify the frequency gap as it changes slightly across the spectrum even though the azimuthal momentum spacing within any mode pair must be the same. We then apply the corresponding 3-phase RF stimulus at that $\Omega$ with $V_{\textrm{RF}}=15.88$ V. This is possible since, as mentioned earlier, the frequency is a completely free variable in this EO design. We observed that the isolator maintains $41-52$ dB contrast and $1.4-2.0$ dB insertion loss for all mode pairs across this spectral range, confirming consistent optical isolation performance. Thus, each mode pair indicates a coarse tuning choice for the isolator.
@@
Finally, we explore the possibility of fine-tuning to fill the spectral gaps between the coarse tuning points. We measure that a dc voltage on the EO tuning electrodes (see Supplementary Data Fig. 3) linearly shifts both modes of a pair together (Fig. 5b). For the device tested the tuning capability was measured at 0.20 GHz/V for the TEodd mode and 0.21 GHz/V for the TEeven mode, with the difference arising due to the distinction of modal group index. The small monotonic change in the frequency gap can be tracked while tuning the device to appropriately adjust the stimulus frequency $\Omega$. Since the free spectral range between optical mode pairs in this device is around $40$ GHz, an application of $\pm$ 100 VDC can tune the modes to completely cover the gaps. We leave further improvements on the tuning structure to future efforts, since the resonator free spectral range can be reduced by simple redesign, and we additionally note that the combination of EO tuning and stress-based piezoelectric tuning [tian_piezoelectric_2024] could provide enhanced tuning capability.
@@ Discussion
This EO isolator delivers $>30$ dB/dB isolator figure-of-merit while using nearly 20x smaller interaction length (900 µm here, vs 1.75 cm) and requires nearly 10x lower RF input power than the best comparable EO isolator (12.5 mW here, vs 125 mW) [yu_integrated_2023]. This is because the stimulus electrodes do not draw RF power from the signal generator, but merely act as open-circuited capacitors. Also, this design consumes nearly 64x lower RF input power than the best comparable performance AO p-ATS isolator of similar design (794 mW) [sohn_electrically_2021]. The efficiency of our device can be improved further by incorporating more periods of the EO stimulus electrodes around the resonator resulting in a larger interaction length. The EO tuning could then be performed using these same stimulus electrodes but with two different dc ground offset voltages. As with the AO p-ATS isolator, our EO p-ATS design has an instantaneous isolation bandwidth that is fundamentally set by the optical modes, though that could be improved somewhat by lowering the optical Q-factors while simultaneously increasing $G_{eo}$. We also showed that very wide tuning capability is enabled thanks to dual innovations on a double-racetrack resonator design and a synthetic momentum approach, both of which together relax the phase-matching constraints. Most importantly, the p-ATS approach intrinsically produces minimal sidebands, but which also improve monotonically as the stimulus level is increased. Thus, future modulation efficiency enhancements with novel materials like BaTiO3 [dong_monolithic_2023, demkov_ferroelectric_2024, mohl_bidirectional_2025] could improve ideality and performance metrics even further. Another approach to reduce the residual sideband is to fully decouple the auxiliary TEeven mode from the waveguide by means of smart optical design. Ultimately, these narrowband but widely tunable non-magnetic isolators can have a big positive impact for single-frequency protection, e.g. for ultrastable laser sources [poulton_coherent_2017, spencer_optical-frequency_2018, lucas_ultralow-noise_2020], atomic references [knappe_microfabricated_2004, hummon_photonic_2018, newman_architecture_2019], and optical metrology [delhaye_optical_2007].
@@ Acknowledgments
This work was sponsored by the Defense Advanced Research Projects Agency (DARPA) under Cooperative Agreement D24AC00003, the Air Force Research Laboratory (AFRL) / US Space Force (USSF) grant FA9453-20-2-0001, and the US Office of Naval Research (ONR) Multi-University Research Initiative grant N00014-20-1-2325. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of DARPA, AFRL, ONR, or the US Government. The authors would like to acknowledge invaluable support from Dr. Brian Kasch at the Air Force Research Laboratory, Space Vehicles Directorate, and Dr. Edmond Chow and Amr O. Ghoname at the Holonyak Micro & Nanotechnology Lab (HMNTL) at University of Illinois at Urbana-Champaign.
@@
Supplementary Data Figure 3: (a) The cross-sectional schematic of the EO tuner on the ring resonator. (b) Example transmission measurements through the waveguide showing how the resonator’s even and odd modes shift as a function of applied dc voltage.
Supplementary Data Table 1: Comparison between published EO and electrically driven AO isolators in integrated photonics.
{adjustwidth}
-1.1in-0.5in Author Year Device Technique Isolation contrast (IC) Insertion loss (IL) IC/IL (dB/dB) This work - Resonator EO 47.7 dB 1.45 dB 32.9 Yu et. al. [yu_integrated_2023] 2023 Waveguide EO 48 dB 1.4 dB (w/ filter) 34.28 0.5 dB (w/o necessary filter) 96 Gao et. al. [gao_thin-film_2024] 2023 Waveguide EO 39.5 dB 2.6 dB (w/o filter) 15.19 Dostart et. al. [dostart_optical_2021] 2021 Resonator EO 13.1 dB 18.1 dB 0.18 Doerr et. al. [doerr_silicon_2014] 2014 Waveguide EO 6 dB 4 dB 0.66 Tzuang et. al. [tzuang_non-reciprocal_2014] 2014 Waveguide EO 2.4 dB N/A N/A Lira et. al. [lira_electrically_2012] 2012 Waveguide EO 3 dB 70 dB 0.04 Doerr et. al. [doerr_optical_2011] 2011 Waveguide EO 2 dB 11 dB 0.18 Bhandare et. al. [bhandare_novel_2005] 2005 Waveguide EO 30 dB 8 dB 3.75 Sohn et. al. [sohn_electrically_2021] 2021 Resonator AO 39.31 dB 0.65 dB 60.46 12.75 dB 1.13 dB 11.28 Tian et. al. [tian_magnetic-free_2021] 2021 Resonator AO 9.3 dB 0.8 dB 11.62 41 dB 1.9 dB 21.57

### v2

@@ Results
[…]
@@
Tunability across a multi-THz spectrum – A wide transmission spectrum of the device, at the limits of our 1520–1570 nm tunable laser, is presented in Fig. 5a. The frequency gap between modes of azimuthal order $\Delta M_{\textrm{optical}}=7$ evolves with a linear trend of $\sim$4.1 GHz/THz across the spectrum due to dispersion. The shape of the background transmission is defined by the optical grating couplers and can be flattened if we instead used edge couplers. We tested the tuning capability of the isolator over the range 193.35–194.45 THz ($\sim$8 nm from 1542–1550 nm) which was only limited by the frequency constraint of our laboratory’s RF signal generators. For all mode pairs in this range, we applied the 3-phase RF stimulus matching the frequency gap $\Omega$ with $V_{\textrm{RF}}=15.88$ V. This is possible since, as mentioned earlier, the frequency is a completely free variable in this EO design. Fig. 5b presents the measured isolation contrast and insertion loss. We confirm that the device has consistent optical isolation performance with 415241-52 dB contrast and 1.42.01.4-2.0 dB insertion loss for all mode pairs across this spectral range. Thus, each mode pair indicates a coarse tuning choice for the isolator.
This coarse tuning capability is not constrained by any fundamental limit, and the isolation can extend far beyond the demonstrated span as long as an RF source is available to bridge the inter-modal frequency gap. The spacing for $\Delta M_{\textrm{optical}}=7$ mode pairs near 191 THz is only $\sim$16 GHz (Fig. 5a). The mode spacing approaches zero near 195 THz, which will be a detriment to isolator directionality, thus our device should have a minimum operational range of 191–194.5 THz. At higher optical frequencies >>195 THz, the modes are again resolved but their spacing is negative which simply means that the isolator will operate in a flipped mode [sohn_time-reversal_2018]. Assuming a continuing linear trend for the dispersions, the mode spacing should reach 60 GHz around at 180 THz optical frequency, which is readily accessible using microwave signal generators, and we can therefore speculate on a potential operational range of 180–194.5 THz (1542–1665 nm). Of course, it is additionally important to modify the device to a broadband optical edge coupling setup to circumvent the bandwidth limitations of the grating couplers which were used in our experiments.
@@
Finally, we explore the possibility of fine-tuning to fill the spectral gaps between the coarse tuning points. We measure that a dc voltage on the co-integrated EO tuning electrodes (see Supplementary Data Fig. 3) linearly shifts both modes of a pair together. For the device tested the tuning capability was measured at 0.20 GHz/V for the TEodd mode and 0.21 GHz/V for the TEeven mode, with the difference arising due to the distinction of modal group index. The small monotonic change in the frequency gap can be tracked while tuning the device to appropriately adjust the stimulus frequency $\Omega$. Since the free spectral range between optical mode pairs in this device is around $40$ GHz, an application of $\pm$ 100 VDC can in principle electro-optically tune the modes to completely cover the gaps. We leave further improvements on the tuning structure to future efforts, since the resonator free spectral range can be reduced by simple redesign, and we additionally note that the combination of EO tuning and stress-based piezoelectric tuning [tian_piezoelectric_2024] could provide enhanced tuning capability. The multi-THz tuning capability, in conjunction with state-of-the-art IC/IL figure of merit, implies unique isolation coverage that far surpasses any previous on-chip isolator.
@@ Discussion
This EO isolator is the first integrated photonic device to demonstrate non-reciprocal strong coupling with electro-optic stimulus. Since isolators can be cascaded, a highly valued figure of merit is the contrast that can be obtained per dB of insertion loss (ratio IC/IL in dB/dB) for the optical carrier signal. Commercial off-chip magneto-optic isolators typically exhibit IC/IL >30>30 dB/dB, a level that to date only two reported non-magnetic chip-scale isolators have reached [sohn_electrically_2021, yu_integrated_2023] (see Supplementary Data Table 1). The device we demonstrate here also reaches this figure-of-merit regime while using nearly 20x smaller linear footprint (900 µm here, vs 1.75 cm) and generating far fewer sidebands than the best comparable EO isolator [yu_integrated_2023]. The efficiency of our device can be improved further by incorporating more periods of the EO stimulus electrodes around the resonator resulting in a larger interaction length. The EO tuning could then be performed using these same stimulus electrodes but with two different dc ground offset voltages. As with the AO chiral shunt isolator, our EO design has an instantaneous isolation bandwidth that is fundamentally set by the optical modes, though that could be improved somewhat by lowering the optical Q-factors while simultaneously increasing $G_{eo}$. Specifically, the maximum isolation bandwidth for our device is limited to the linewidth of the TEodd mode, even for infinite $G\textsubscript{eo}$. We also showed that very wide tuning capability is enabled thanks to dual innovations on a double-racetrack resonator design and a synthetic momentum approach, both of which together relax the phase-matching constraints. Thus, future modulation efficiency enhancements with novel materials like BaTiO3 [dong_monolithic_2023, demkov_ferroelectric_2024, mohl_bidirectional_2025], which can now reach the same required Q-factors with much higher Pockels coefficient [kim_low_2025], could improve ideality and performance metrics even further. Another approach to reduce the residual sideband is to fully decouple the auxiliary TEeven mode from the waveguide by means of smart optical design. Ultimately, these narrowband but widely tunable non-magnetic isolators can have a big positive impact for single-frequency protection, e.g. for ultrastable laser sources [poulton_coherent_2017, spencer_optical-frequency_2018, lucas_ultralow-noise_2020], atomic references [knappe_microfabricated_2004, hummon_photonic_2018, newman_architecture_2019], and optical metrology [delhaye_optical_2007].
@@ Acknowledgments
This work was sponsored by the Defense Advanced Research Projects Agency (DARPA) under Cooperative Agreement D24AC00003, the Air Force Research Laboratory (AFRL) / US Space Force (USSF) grant FA9453-20-2-0001, the US Office of Naval Research (ONR) Multi-University Research Initiative grant N00014-20-1-2325, and the Army Research Office (ARO) grant W911NF-23-1-0219. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of DARPA, AFRL, ONR, ARO, or the US Government. The authors would like to acknowledge invaluable support from Dr. Brian Kasch at the Air Force Research Laboratory, Space Vehicles Directorate, and Dr. Edmond Chow and Amr O. Ghoname at the Holonyak Micro & Nanotechnology Lab (HMNTL) at University of Illinois at Urbana-Champaign.
Data availability
The data that support the findings of this study are available from the corresponding authors upon reasonable request.
Author contributions
G.I.K. and G.B. jointly conceived the concept, G.I.K. and V.W. performed the theoretical analysis. G.I.K., O.E.Ö., and J.Y. contributed to the device fabrication and the experimental measurements. G.I.K. and V.W. analyzed the experimental data. All authors contributed to writing the paper. G.B. supervised all aspects of this project.
@@
Supplementary Data Figure 3: (a) The cross-sectional schematic of the EO tuner on the ring resonator (shown in Fig. 2e). (b) Example transmission measurements through the waveguide, on a device with mode pair spacing around 11.5 GHz, showing how the resonator’s even and odd modes shift with applied dc voltage. (c) Consolidated tracking of the individual mode detuning and evolution of the frequency gap as a function of applied dc voltage.
Supplementary Data Table 1: Comparison between published EO and electrically driven AO isolators in integrated photonics. Results with IC/IL >30>30 dB/dB, i.e. competitive with commercial isolators, are marked in bold.
Author Year Device Technique Isolation contrast (IC) Insertion loss (IL) IC/IL (dB/dB)
[…]